Googl ngphysi PHYS 16 Problem 6.6 Problem BA In ea
Answer #1

Part\:A:

Applying\:Second\:Newton's\:Law:

\sum \vec{F}=m\vec{a}

To\:obtain\:acceleration\:on\:x-components,\:we\:need\:forces\:on\:same

components.\:The\:values\:of\:the\:larger\:vector,\:are\:its\:components.

F_x=ma_x

2\:N-4\:N=\left ( 2\:kg \right )a_x

solving\:for\:a_x:

a_x=\frac{2\:N-4\:N}{2\:kg}={\color{Blue} -1\:\frac{m}{s^{2}}}

Part B

Similar\:way\:from\:Part\:A:

F_y=ma_y

a_y=\frac{F_y}{m}=\frac{3\:N-3\:N}{2\:kg}={\color{Blue} 0\:\frac{m}{s^{2}}}

{\color{Red} NOTE:}\:For\:Parts\:C\:and\:D,\:you\:should\:provide\:the\:figure\:2.

However,\:you\:can\:follow\:steps\:done\:in\:Part\:A\:and\:B.